An architect designs a rectangular flower garden such that the width is exactly 2/3 of the length. If 380 feet of antique Picket fencing are to be used to enclose the garden, find the dimensions of the garden
Answer by addingup(3677)

Width = 2/3 Length
and
2L+2W = Perimeter Substitute with what we know:
2L+2(2/3L) = 380
2L+4/3L = 380 Put both numbers on the left over a common denominator
6/3L+4/3L = 380
10/3L = 380 divide both sides by 10/3
L = 380*3/10 = 114 is the length, and the width:
W = 2/3L = 2/3*114 = (2*114)/3 = 76 is the width
>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>
Check:
2L+2W = Perimeter
2(114)+2(76) = 380
228+152 = 380 Correct (add the left, you get 380 as you should)
:
John